OpenStudy (aleah54):

FAN AND MEDAL The lengths of two sides of a triangle are shown below: Side 1: 3x2 - 4x - 1 Side 2: 4x - x2 + 5 The perimeter of the triangle is 5x3 - 2x2 + 3x - 8. Part A: What is the total length of the two sides, 1 and 2, of the triangle? (4 points) Part B: What is the length of the third side of the triangle? (4 points) Part C: Do the answers for Part A and Part B show that the polynomials are closed under addition and subtraction? Justify your answer

OpenStudy (aleah54):

I got a and b

OpenStudy (aleah54):

@dancergirl45 @Datman21328 @Destinyyyy @Daniellelovee @danielmanzo @FlakeSnow @fluttershyK

OpenStudy (aleah54):

@Zeronknight @zaberangelr @zainabkhan

OpenStudy (aleah54):

@jigglypuff314 @Ashleyisakitty

OpenStudy (aleah54):

@imqwerty @chainedecho @cruffo @cdobbs2017 @ChaosGuardian14

OpenStudy (aleah54):

@jhonyy9

jhonyy9 (jhonyy9):

just you ned solve these quadratics - do you can ?

jhonyy9 (jhonyy9):

@toriacm18 sorry but i not can seniding messages bc. my submit button is blocked again too

jhonyy9 (jhonyy9):

3x^2 -4x -1 =0 this is the one from these two quadratics so how is this second ?

OpenStudy (aleah54):

here is my answer for a and b Part A Total length = side 1 + side 2 8x^2 - 5x - 2 + 7x - x^2 + 3 so the answer is 7x^2 + 2x + 1 Part B perimeter = a + b + c 4x^3- 3x^2 + 2x - 6 = 8x^2 - 5x - 2 + 7x - x^2 + 3 + c 4x^3 - 3x^2 + 2x - 6 = 7x^2 + 2x + 1 + c 4x^3 - 3x^2 - 7x^2 + 2x - 2x - 1 - 6 so the answer is 4x^3 - 10x^2 - 7

OpenStudy (aleah54):

I still don't get part c

OpenStudy (aleah54):

@jhonyy9

OpenStudy (aleah54):

Can someone please help

OpenStudy (aleah54):

i got it